# Efficient Adaptive Experimentation with Non-Compliance

### v1

@@ A.2 Auxiliary Context
Appendix B Asymptotic Confidence Sequences
@@
[…]
Constructing non-asymptotic, anytime-valid CSs can be difficult in when the target estimand contains estimated nuisance functions. Fortunately, for AMRIV the nuisance-induced remainder is $o_{p}(t^{-1/2})$ under Theorem 3 assumptions, so an asymptotic CS—valid after a finite, burn-in phase—in remains both tractable and practically useful.
@@
[…]
@@
Suppose Assumptions 1, 2 and 3 hold and there exists a non-adaptive policy $\pi(X)\in[\epsilon,1-\epsilon]$ for some $\epsilon>0$ such that the nuisances estimates $\widehat{\eta}_{t}$ and the adaptive assignment policy $\pi_{t}(X\mid\mathcal{H}_{t-1})$ are $L_{2}$-consistent relative to the truncation schedule, i.e. $k_{t}\|\widehat{f}_{t-1}-f\|_{2}=o_{p}(1)$ and $k_{t}\|\pi_{t}-\pi\|_{2}=o_{p}(1)$ for $f\in\{\mu^{Y}(0,\cdot),\mu^{A}(0,\cdot),\delta(\cdot),\delta^{A}(\cdot)\}$. Furthermore, assume $\|\hat{\delta}_{t-1}-\delta\|_{2}=o_{a.s.}(1)$ and $\|\hat{\delta}_{t-1}-\delta\|_{2}\|\widehat{\delta}_{t-1}^{A}-\delta^{A}\|_{2}=o_{a.s.}\left(\sqrt{\frac{\log t}{t}}\right)$. Let
@@
$|\phi(X,Z,A,Y)|$ $=\frac{2}{\epsilon}\left(3C^{2}+C^{3}\right)+C.$
@@
Because the constant $C$ is finite, the variance must also be finite, which completes our proof.
By Lemma 7, the confidence sequence in Theorem 6 is a valid $(1-\alpha)$-AsympCS for $\tau$. The proof for the approximate choice of $\rho^{*}$ that minimizes the relative width of the confidence sequence at time $T$ is provided in Waudby-Smith et al. [45, Appendix B.2].
Appendix C Proof of Theorem 1 and Corollary 2
@@
$=\mathbb{E}\left[\mathbb{E}[(\phi(X,Z,A,Y;\eta)-\tau)^{2}\mid X]\right]$ (Law of iteratred expectations)
@@
$\quad=\underbrace{\frac{2Z-1}{Z\pi(X)+(1-Z)(1-\pi(X))}\frac{1}{\delta^{A}(X)}\left[Y-A\delta(X)-\mu^{Y}(0,X)+\mu^{A}(0,X)\delta(X)\right]}_{\Lambda_{1}}+\delta(X)-\tau$
@@
In now remains to expand $\mathbb{E}[\Lambda_{1}^{2}\mid X]$:
@@
$=\frac{1}{\delta^{A}(X)^{2}}\left(\frac{\sigma^{2}(1,X)}{\pi(X)}\frac{\sigma^{2}(0,X)}{1-\pi(X)}\right)$
@@
The minimum is obtained when the derivative of the argument w.r.t. $p$ is 00, i.e. $\frac{\sigma^{2}(0,X)}{(1-p)^{2}}-\frac{\sigma^{2}(1,X)}{p^{2}}=0$. By solving for $p$, we obtain $p=\frac{\sqrt{\sigma^{2}(1,X)}}{\sqrt{\sigma^{2}(1,X)}+\sqrt{\sigma^{2}(0,X)}}$. Thus, we obtain the result of Corollary 2:
@@
Appendix D Proof of Theorem 3
D.1 Preliminaries
Our asymptotic argument relies on a martingale central limit theorem under a Lindeberg-type condition. We use a streamlined version of the MDS central limit theorem originally due to Dvoretzky [17], as presented in Zhang et al. [49, Theorem 2].
Theorem 8 (Martingale CLT, adapted from [49, Thm. 2]).
@@ D.1 Preliminaries
[…]
@@
[…]
@@
[…]
To begin our proof, we define $\psi_{t}:=\phi(X_{t},A_{t},Z_{t},Y_{t};\pi_{t},\widehat{\eta}_{t})-\tau$, where $\phi$ is given in Eq. (5.2). Letting $\eta=\{\mu^{Y}_{t-1}(0,X),\mu^{A}_{t-1}(0,X),\delta_{t-1}^{A}(X),\delta_{t-1}(X)\}$ denote the true nuisance values, we decompose $\psi_{t}$ as
@@
[…]
D.2 MDS structure of $z_{t}$
@@
D.3 $z_{t}$ satisfies conditions (2)–(3) of Theorem 8
[…]
@@ D.3 $z_{t}$ satisfies conditions (2)–(3) of Theorem 8
[…]
@@
[…]
@@
[…]
@@
[…]
@@
D.4 $\sqrt{T}\left(\frac{1}{T}\sum_{t=1}^{T}m_{t}\right)$ is $o_{p}(1)$
@@
$\qquad\Bigg{\{}Y_{t}\frac{\delta^{A}(X_{t})-\widehat{\delta}^{A}_{t-1}(X_{t})}{\delta^{A}(X_{t})\widehat{\delta}^{A}_{t-1}(X_{t})}-A_{t}\left(\frac{\widehat{\delta}_{t-1}(X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\delta(X_{t})}{\delta^{A}(X_{t})}\right)+\left(\frac{\widehat{\mu}^{Y}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\mu^{Y}(0,X_{t})}{\delta^{A}(X_{t})}\right)$
$\qquad\quad+\left(\frac{\widehat{\mu}^{A}_{t-1}(0,X_{t})\widehat{\delta}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\mu^{A}_{t-1}(0,X_{t})\delta_{t-1}(0,X_{t})}{\delta^{A}(X_{t})}\right)+(\widehat{\delta}(X_{t})-\delta(X_{t}))\Bigg{\}}.$
@@
$\quad+2\mathbb{E}\left[\frac{1}{\pi_{t}(X_{t})}\mathbb{E}[A_{t}^{2}\mid Z_{t}=1,X_{t}]\left(\frac{\widehat{\delta}_{t-1}(X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\delta(X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr{|}\mathcal{H}_{t-1}\right]$ (b)
$\quad+2\mathbb{E}\left[\frac{1}{\pi_{t}(X_{t})}\left(\frac{\widehat{\mu}^{Y}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\mu^{Y}(0,X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr{|}\mathcal{H}_{t-1}\right]$ (c)
$\quad+2\mathbb{E}\left[\frac{1}{\pi_{t}(X_{t})}\left(\frac{\widehat{\mu}^{A}_{t-1}(0,X_{t})\widehat{\delta}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\mu^{A}_{t-1}(0,X_{t})\delta_{t-1}(0,X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr{|}\mathcal{H}_{t-1}\right]$ (d)
$\quad+2\mathbb{E}\left[(\widehat{\delta}(X_{t})-\delta(X_{t}))^{2}\big{|}\mathcal{H}_{t-1}\right]$ (e)
@@
$\quad-2\mathbb{E}\left[\frac{1}{1-\pi_{t}(X_{t})}\mathbb{E}[A_{t}^{2}\mid Z_{t}=1,X_{t}]\left(\frac{\widehat{\delta}_{t-1}(X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\delta(X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr{|}\mathcal{H}_{t-1}\right]$ (b)
$\quad-2\mathbb{E}\left[\frac{1}{1-\pi_{t}(X_{t})}\left(\frac{\widehat{\mu}^{Y}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\mu^{Y}(0,X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr{|}\mathcal{H}_{t-1}\right]$ (c)
$\quad-2\mathbb{E}\left[\frac{1}{1-\pi_{t}(X_{t})}\left(\frac{\widehat{\mu}^{A}_{t-1}(0,X_{t})\widehat{\delta}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}(X_{t})}-\frac{\mu^{A}_{t-1}(0,X_{t})\delta_{t-1}(0,X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr{|}\mathcal{H}_{t-1}\right]$ (d)
$\quad-2\mathbb{E}\left[(\widehat{\delta}(X_{t})-\delta(X_{t}))^{2}\big{|}\mathcal{H}_{t-1}\right]$ (e)
@@
$\leq 4k_{t}C^{4}\epsilon_{\delta^{A}}^{2}\|\widehat{\delta}^{A}-\delta^{A}\|_{2}^{2}$ (a)
@@
$\quad+\|\widehat{\delta}(X_{t})-\delta(X_{t})\|_{2}^{2}$ (e)

### v2

@@ A.2 Auxiliary Context
Appendix B Notation
Table 1: Notation
$X$ Observed covariates (feature vector) in $\mathbb{R}^{m}$.
$Z$ Binary instrument / encouragement, $Z\in\{0,1\}$ (experimenter-assigned).
AA Binary treatment, $A\in\{0,1\}$, determined endogenously by the unit.
$Y$ Real-valued observed outcome.
$Y(a)$, $Y(a,z)$ Potential outcome under treatment $a$ (and instrument $z$ when shown).
$A(z)$ Potential treatment taken under instrument $z$.
$U$ Unobserved confounder(s).
$P$ True (observable) distribution of $(X,Z,A,Y)$.
$\tau$ Population average treatment effect (ATE): $\tau:=\mathbb{E}[Y(1)]-\mathbb{E}[Y(0)]$.
$\mu^{Y}(z,x)$ $\mathbb{E}[Y\mid Z=z,X=x]$, instrument-conditional outcome regression.
$\mu^{A}(z,x)$ $\mathbb{E}[A\mid Z=z,X=x]$, instrument-conditional treatment regression.
$\delta^{Y}(x),\,\delta^{A}(x)$ Instrument-induced shifts in $Y$ and $A$, respectively.
$\delta^{Y}(x)=\mu^{Y}(1,x)-\mu^{Y}(0,x)$, $\delta^{A}(x)=\mu^{A}(1,x)-\mu^{A}(0,x)$:
$\delta(x)$ Conditional average treatment effect (CATE) at $x$: $\delta(x):=\delta^{Y}(x)/\delta^{A}(x)$.
$\pi_{t}(x\mid\mathcal{H}_{t-1})$ Instrument assignment policy at time $t$: $\Pr(Z_{t}=1\mid X_{t}=x,\mathcal{H}_{t-1})$.
$\pi^{*}(x)$ Efficiency-optimal instrument assignment (Eq. (4)).
$\mathcal{H}_{t-1}$ History before round $t$: $\{(X_{i},Z_{i},A_{i},Y_{i})\}_{i=1}^{t-1}$.
$T$ Total number of rounds / samples.
$T_{0}$ Burn-in rounds (initial exploration).
$k_{t}$ Truncation parameter at time $t$ (ensures $\pi_{t}\in[1/k_{t},1-1/k_{t}]$).
$\sigma^{2}(z,x)$ Residual variance: $\operatorname*{{\rm Var}}(Y-A\delta(X)\mid Z=z,X=x)$.
$V_{\mathrm{eff}}(\pi)$ Semiparametric efficiency bound under policy $\pi$ (Eq. (3)).
$\phi(\cdot;\pi,\eta)$ Recentered efficient influence function (EIF) used in AMRIV (Eq. (2)).
$\widehat{f}$, $\widetilde{f}$ $\widehat{f}$: estimate of $f$; $\widetilde{f}$: plug-in or candidate function.
$\widehat{\tau}^{\mathrm{AMRIV}}_{T}$ AMRIV estimator after $T$ rounds (Alg. 1).
$\widehat{\mathbb{E}}$ Empirical expectation (sample average).
$\|g\|_{2}$ $L_{2}$ norm: $\|g\|_{2}:=\mathbb{E}[g(X)^{2}]^{1/2}$.
$o_{p}(1)$, $O_{p}(\cdot)$ Standard probabilistic asymptotic order notation.
$\varepsilon$ Small positive constant (e.g., variance-floor in Eq. (6)).
$\mathcal{H}_{t-1}^{(j)}$ Temporal cross-fitting fold (e.g., $j\in\{0,1\}$) used for sequential cross-fitting.
Appendix C Practical Implementation of Nuisance and Variance Estimators
This section complements Remark 2 by outlining practical choices for nuisance and variance estimators, including nonnegativity and online-update considerations.
Nuisance estimators.
Because AMRIV uses sequential cross-fitting, we require only standard nonparametric convergence rates to establish asymptotic normality (Theorems 34). Consequently, the analyst may flexibly choose estimators according to data structure and sample size. Representative options include:
Table 2: Representative nuisance estimators, convergence rates, and suitable applications.
Estimator Convergence rate Best suited for
$k$-NN / kernel smoother $O\!\left(n^{-\beta/(2\beta+d)}\right)$ for Hölder-$\beta$ smooth functions [41] Low-dimensional, smooth problems
Random forest $\tilde{O}\!\left(n^{-\beta/(2\beta+d)}\right)$ for Hölder-$\beta$ smooth functions [49, 53] Moderate-dimensional, tabular data
Fully connected neural nets (ReLU) $O\!\left(\sqrt{WL\log W}\,n^{-1/2}\right)$ for width $W$ and depth $L$ [59] High-dimensional or structured inputs
Neural nets (1-Lipschitz, bounded weights) $O\!\left(\sqrt{\prod_{l=1}^{L}M_{l}}/n^{1/4}\right)$ where $M_{l}$ bounds the Frobenius norm of layer $l$’s weights [20] High-dimensional or structured inputs
The best choice depends on sample size, covariate dimension, and smoothness. In adaptive experiments with directly assigned treatments, $k$-NN and random forests are standard baselines [27, 13].
Online or streaming implementations.
Theoretical results assume that nuisance functions are re-estimated at each update, but full data storage is not required. Practical alternatives include:
  1. 1.
    Online learners: gradient-descent (SGD) updates, online random forests [32], or GLMs can update parameters incrementally using new data.
  2. 2.
    Block sample-splitting: reuse recent batches to update nuisance fits and discard older data while maintaining cross-fitting validity [26].
  3. 3.
    Sufficient-statistic updates: for parametric or binned regressors, sufficient statistics such as running sums of $(X,Z,A,Y)$ per cell suffice.
Non-negative variance estimation.
Equation (6) enforces non-negativity via a small floor parameter $\varepsilon$. As an alternative, one may use a self-normalized kernel variance estimator that is fully nonparametric and guarantees $\widehat{\sigma}^{2}_{t}(z,x)\geq 0$ by construction.
For each $(z,x)$, define
$\widehat{\sigma}^{2}_{t}(z,x)$ $=\frac{\sum_{s\leq t}K_{h}(\|X_{s}-x\|)\,\mathbb{I}\{Z_{s}=z\}\,(R_{s}-\widehat{\mu}_{t-1}(z,x))^{2}}{\sum_{s\leq t}K_{h}(\|X_{s}-x\|)\,\mathbb{I}\{Z_{s}=z\}},$
$\widehat{\mu}_{t-1}(z,x)$ $=\frac{\sum_{s\leq t}K_{h}(\|X_{s}-x\|)\,\mathbb{I}\{Z_{s}=z\}\,R_{s}}{\sum_{s\leq t}K_{h}(\|X_{s}-x\|)\,\mathbb{I}\{Z_{s}=z\}},$
where $R_{s}=Y_{s}-A_{s}\,\widehat{\delta}_{t-1}(X_{s})$ and $K_{h}(\cdot)$ is a kernel with bandwidth $h$.
This estimator satisfies $\widehat{\sigma}^{2}_{t}(z,x)\geq 0$ by construction and is consistent for $\operatorname*{{\rm Var}}(Y-A\delta(X)\mid Z=z,X=x)$ under standard kernel conditions. Caveat: updating kernel weights online can be $O(t^{2})$ in general; for large-scale data, the plug-in estimator in Eq. (6) is more computationally efficient.
Both the plug-in estimator with the floor $\varepsilon$ and the kernel-based version above preserve all asymptotic guarantees; the choice between them is primarily a trade-off between computational efficiency and strict non-negativity.
Appendix D Asymptotic Confidence Sequences
@@
$P\left(\,\forall t\in\mathbb{N}^{+}:\tau\in[L_{t},U_{t}]\right)\;\geq\;1-\alpha.$
Constructing non-asymptotic, anytime-valid CSs can be difficult in when the target estimand contains estimated nuisance functions. Fortunately, for AMRIV the nuisance-induced remainder is $o_{p}(t^{-1/2})$ under Theorem 3 assumptions, so an asymptotic CS—valid after a finite, burn-in phase— remains both tractable and practically useful.
@@
    $\lim_{T_{0}\to\infty}P\!\left(\forall t\geq T_{0}:\,\tau\in\widetilde{C}_{t}\right)\;\geq\;1-\alpha.$
@@
Suppose Assumptions˜1, 2 and 3 hold and there exists a non-adaptive policy $\pi(X)\in[\epsilon,1-\epsilon]$ for some $\epsilon>0$ such that the nuisances estimates $\widehat{\eta}_{t}$ and the adaptive assignment policy $\pi_{t}(X\mid\mathcal{H}_{t-1})$ are $L_{2}$-consistent relative to the truncation schedule, i.e. $k_{t}\|\widehat{f}_{t-1}-f\|_{2}=o_{p}(1)$ and $k_{t}\|\pi_{t}-\pi\|_{2}=o_{p}(1)$ for $f\in\{\mu^{Y}(0,\cdot),\mu^{A}(0,\cdot),\delta(\cdot),\delta^{A}(\cdot)\}$. Furthermore, assume $\|\hat{\delta}_{t-1}-\delta\|_{2}=o_{a.s.}(1)$ and $\|\hat{\delta}_{t-1}-\delta\|_{2}\|\widehat{\delta}_{t-1}^{A}-\delta^{A}\|_{2}=o_{a.s.}\left(\sqrt{\frac{\log t}{t}}\right)$. Let
@@
$|\phi(X,Z,A,Y;\pi,\eta)|$ $=\frac{2}{\epsilon}\left(3C^{2}+C^{3}\right)+C.$
@@
Because the constant $C$ is finite, the variance must also be finite, which completes our proof. By Lemma 7, the confidence sequence in Theorem 6 is a valid $(1-\alpha)$-AsympCS for $\tau$. The proof for the approximate choice of $\rho^{*}$ that minimizes the relative width of the confidence sequence at time $T$ is provided in Waudby-Smith et al. [55, Appendix B.2]. ∎
Appendix E Proof of Theorem 1 and Corollary 2
@@
$=\mathbb{E}\left[\mathbb{E}[(\phi(X,Z,A,Y;\eta)-\tau)^{2}\mid X]\right]$ (Law of iterated expectations)
@@
$\qquad=\underbrace{\frac{2Z-1}{Z\pi(X)+(1-Z)(1-\pi(X))}\frac{1}{\delta^{A}(X)}\left[Y-A\delta(X)-\mu^{Y}(0,X)+\mu^{A}(0,X)\delta(X)\right]}_{\Lambda_{1}}+\delta(X)-\tau$
@@
It now remains to expand $\mathbb{E}[\Lambda_{1}^{2}\mid X]$:
@@
$=\frac{1}{\delta^{A}(X)^{2}}\left(\frac{\sigma^{2}(1,X)}{\pi(X)}+\frac{\sigma^{2}(0,X)}{1-\pi(X)}\right)$
@@
The minimum is obtained when the derivative of the argument w.r.t. $p$ is 0, i.e. $\frac{\sigma^{2}(0,X)}{(1-p)^{2}}-\frac{\sigma^{2}(1,X)}{p^{2}}=0$. By solving for $p$, we obtain $p=\frac{\sqrt{\sigma^{2}(1,X)}}{\sqrt{\sigma^{2}(1,X)}+\sqrt{\sigma^{2}(0,X)}}$. Thus, we obtain the result of Corollary 2:
@@
Appendix F Proof of Theorem 3
F.1 Preliminaries
Our asymptotic argument relies on a martingale central limit theorem under a Lindeberg-type condition. We use a streamlined version of the MDS central limit theorem originally due to Dvoretzky [17], as presented in Zhang et al. [60, Theorem 2].
Theorem 8 (Martingale CLT, adapted from [60, Thm. 2]).
@@ F.1 Preliminaries
    $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[z_{t}^{2}\mid\mathcal{H}_{t-1}]\overset{p}{\longrightarrow}\sigma^{2},$
@@
    $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[z_{t}^{2}\mathbb{I}\{|z_{t}|>\epsilon\sqrt{T}\}\mid\mathcal{H}_{t-1}\right]\overset{p}{\longrightarrow}0.$
@@
$\sqrt{T}\;\overline{z}_{T}\overset{d}{\longrightarrow}\mathcal{N}(0,\sigma^{2}).$
To begin our proof, we define $\psi_{t}:=\phi(X_{t},A_{t},Z_{t},Y_{t};\pi_{t},\widehat{\eta}_{t})-\tau$, where $\phi$ is given in Eq. (5.2). Letting $\eta=\{\mu^{Y}(0,X),\mu^{A}(0,X),\delta^{A}(X),\delta(X)\}$ denote the true nuisance values, we decompose $\psi_{t}$ as
@@
$\sqrt{T}(\widehat{\tau}^{\text{AMRIV}}_{T}-\tau)\overset{d}{\longrightarrow}\mathcal{N}(0,V_{\text{eff}}(\pi)).$
F.2 MDS structure of $z_{t}$
@@
F.3 $z_{t}$ satisfies conditions (2)–(3) of Theorem 8
For condition (2), we first show that $\mathbb{E}[z_{t}^{2}\mid\mathcal{H}_{t-1}]-V_{\mathrm{eff}}(\pi)\overset{p}{\longrightarrow}0$:
@@ F.3 $z_{t}$ satisfies conditions (2)–(3) of Theorem 8
where in the last line we use the following boundedness conditions: (i) $|Y|\leq C$ from Assumption 3 and thus $|\delta(X)|\leq 2C$ and $\sigma^{2}(z,X_{t})=\mathbb{E}[(Y-A\delta(X_{t}))^{2}\mid Z=z,X_{t}]-\mathbb{E}[Y-A\delta(X_{t})\mid Z=z,X_{t}]^{2}\leq 18C^{2}$, (ii) $|\delta^{A}(X)|^{-1}\leq\epsilon_{\delta^{A}}$ for some $\epsilon_{\delta^{A}}>0$ implicit in the (conditional) relevance in Assumption 1, (iii) $\pi(X_{t}),1-\pi(X_{t})>\epsilon$ from Theorem 3 statement, (iv) $\pi(X_{t}),1-\pi(X_{t})\geq 1/k_{t}$ by construction, and (v) the $L_{1}$ norm is bounded by the $L_{2}$ norm. Thus, setting $\sigma^{2}:=V_{\mathrm{eff}}$, we have that each term converges in probability to $\sigma^{2}$, i.e. $\mathbb{E}[z_{t}^{2}\mid\mathcal{H}_{t-1}]\overset{p}{\longrightarrow}\sigma^{2}$, where $\sigma^{2}$ is finite by Assumption 1 and Assumption 3.
@@
$\left|\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[z_{t}^{2}\mid\mathcal{H}_{t-1}]-\sigma^{2}\right|\overset{p}{\longrightarrow}0.$
Let $a_{t}:=\mathbb{E}[z_{t}^{2}\mid\mathcal{H}_{t-1}]$ and $a:=\sigma^{2}$. We have just established that $a_{t}\overset{p}{\to}a$, and under our boundedness assumptions, $\sup_{t}\mathbb{E}[a_{t}]<\infty$, so the sequence $\{a_{t}\}$ is uniformly integrable. By the $L^{1}$ convergence theorem (e.g., Loève [36]), this implies that $a_{t}\to a$ in $L^{1}$, i.e., $\mathbb{E}\left[\left|\mathbb{E}[z_{t}^{2}\mid\mathcal{H}_{t-1}]-\sigma^{2}\right|\right]\to 0$. Therefore, by Cesàro averaging and Markov’s inequality, we obtain $\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[z_{t}^{2}\mid\mathcal{H}_{t-1}]\overset{p}{\longrightarrow}\sigma^{2}$, completing the verification of condition (2) in Theorem 8.
@@
Since $\operatorname*{{\rm Var}}(z_{t})=\mathbb{E}[z_{t}^{2}]<\infty$, it follows that $\lim_{T\to\infty}\frac{\operatorname*{{\rm Var}}(z_{t})}{\delta^{2}T}=0$, which implies $b_{t}\overset{p}{\to}0$ and hence $b_{t}\overset{d}{\to}0$. Moreover, note that $|b_{t}|\leq z_{t}^{2}$ and $\mathbb{E}[z_{t}^{2}]<\infty$. By the dominated convergence theorem,
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[z_{t}^{2}\cdot\mathbb{I}\left\{|z_{t}|>\delta\sqrt{T}\right\}\mid\mathcal{H}_{t-1}\right]\overset{p}{\longrightarrow}0,$
@@
F.4 $\sqrt{T}\left(\frac{1}{T}\sum_{t=1}^{T}m_{t}\right)$ is $o_{p}(1)$
@@
$\qquad\quad+\widehat{\delta}_{t-1}(X_{t})-\delta(X_{t})\biggr|\mathcal{H}_{t-1}\Bigg]$
$=\mathbb{E}\left[\frac{1}{\widehat{\delta}^{A}_{t-1}(X_{t})}(\delta^{Y}(X_{t})-\delta^{A}(X_{t})\widehat{\delta}_{t-1}(X_{t}))+\widehat{\delta}_{t-1}(X_{t})-\delta(X_{t})\biggr|\mathcal{H}_{t-1}\right]$
$=\mathbb{E}\left[\frac{\delta^{A}(X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}(\delta(X_{t})-\widehat{\delta}_{t-1}(X_{t}))+\widehat{\delta}_{t-1}(X_{t})-\delta(X_{t})\biggr|\mathcal{H}_{t-1}\right]$
@@
$\qquad\Bigg\{Y_{t}\frac{\delta^{A}(X_{t})-\widehat{\delta}^{A}_{t-1}(X_{t})}{\delta^{A}(X_{t})\widehat{\delta}^{A}_{t-1}(X_{t})}-A_{t}\left(\frac{\widehat{\delta}_{t-1}(X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\delta(X_{t})}{\delta^{A}(X_{t})}\right)+\left(\frac{\widehat{\mu}^{Y}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\mu^{Y}(0,X_{t})}{\delta^{A}(X_{t})}\right)$
$\qquad\quad+\left(\frac{\widehat{\mu}^{A}_{t-1}(0,X_{t})\widehat{\delta}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\mu^{A}(0,X_{t})\delta(0,X_{t})}{\delta^{A}(X_{t})}\right)+(\widehat{\delta}_{t-1}(X_{t})-\delta(X_{t}))\Bigg\}.$
@@
$\quad+2\mathbb{E}\left[\frac{1}{\pi_{t}(X_{t})}\mathbb{E}[A_{t}^{2}\mid Z_{t}=1,X_{t}]\left(\frac{\widehat{\delta}_{t-1}(X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\delta(X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr|\mathcal{H}_{t-1}\right]$ (b)
$\quad+2\mathbb{E}\left[\frac{1}{\pi_{t}(X_{t})}\left(\frac{\widehat{\mu}^{Y}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\mu^{Y}(0,X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr|\mathcal{H}_{t-1}\right]$ (c)
$\quad+2\mathbb{E}\left[\frac{1}{\pi_{t}(X_{t})}\left(\frac{\widehat{\mu}^{A}_{t-1}(0,X_{t})\widehat{\delta}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\mu^{A}_{t-1}(0,X_{t})\delta_{t-1}(0,X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr|\mathcal{H}_{t-1}\right]$ (d)
$\quad+2\mathbb{E}\left[(\widehat{\delta}_{t-1}(X_{t})-\delta(X_{t}))^{2}\big|\mathcal{H}_{t-1}\right]$ (e)
@@
$\quad-2\mathbb{E}\left[\frac{1}{1-\pi_{t}(X_{t})}\mathbb{E}[A_{t}^{2}\mid Z_{t}=1,X_{t}]\left(\frac{\widehat{\delta}_{t-1}(X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\delta(X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr|\mathcal{H}_{t-1}\right]$ (b)
$\quad-2\mathbb{E}\left[\frac{1}{1-\pi_{t}(X_{t})}\left(\frac{\widehat{\mu}^{Y}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\mu^{Y}(0,X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr|\mathcal{H}_{t-1}\right]$ (c)
$\quad-2\mathbb{E}\left[\frac{1}{1-\pi_{t}(X_{t})}\left(\frac{\widehat{\mu}^{A}_{t-1}(0,X_{t})\widehat{\delta}_{t-1}(0,X_{t})}{\widehat{\delta}^{A}_{t-1}(X_{t})}-\frac{\mu^{A}_{t-1}(0,X_{t})\delta_{t-1}(0,X_{t})}{\delta^{A}(X_{t})}\right)^{2}\Biggr|\mathcal{H}_{t-1}\right]$ (d)
$\quad-2\mathbb{E}\left[(\widehat{\delta}_{t-1}(X_{t})-\delta(X_{t}))^{2}\big|\mathcal{H}_{t-1}\right]$ (e)
@@
$\leq 4k_{t}C^{4}\epsilon_{\delta^{A}}^{2}\|\widehat{\delta}^{A}_{t-1}-\delta^{A}\|_{2}^{2}$ (a)
@@
$\quad+\|\widehat{\delta}_{t-1}(X_{t})-\delta(X_{t})\|_{2}^{2}$ (e)